OpenStudy (anonymous):

please please help ~ word problem for algebra 1 One printing machine works twice as fast as another. When both machines are used, they can print a magazine in 3 hrs. How many hours would each machine require to do the job alone? thanks in advance xx

OpenStudy (anonymous):

the equation is work rate x time = work done and then work done by A + work done by B = total work done i'm just kinda confused on how to go about setting it up ;3

OpenStudy (mathmale):

Let's refer to the task of printing this magazine as '1 job.' Let the rate at which one machine can do the job be \[\frac{ 1~job }{ x~hrs }~and ~the~other~rate~\frac{ 1~job }{2x~hrs }\] Is this enuf info so that you can do the rest of the setup?

OpenStudy (mathmale):

Find x, and then find 2x, Be certain to check your results.

OpenStudy (anonymous):

so would it be... \[\frac{ 2}{ 2x} + \frac{ 1 }{ 2x} = 1\]

OpenStudy (phi):

you would add the rates of each 1/2x + 1/x (the x is the unknown # of hours) and then you would multiply by the time it takes the two of them to finish 1 job (3 hours according to the problem) so \[ \left(\frac{1}{2x} + \frac{1}{x}\right)3 = 1 \]

OpenStudy (phi):

your equation is almost correct, except you should multiply the left side by 3

OpenStudy (anonymous):

ohh that makes sense :) so using that equation... would the answer be 9?

OpenStudy (phi):

there are 2 answers: 1 for each machine

OpenStudy (anonymous):

i got 9 hours for the first machine and 18 for the second

OpenStudy (phi):

\[ \left(\frac{1}{2x} + \frac{1}{x}\right)3 = 1 \\ \left(\frac{1}{2x} + \frac{2}{2x}\right)3 = 1\] now add the two fractions on the left side. what do you get ?

OpenStudy (anonymous):

\[\frac{ 3 }{ 2x} + \frac{ 6 }{ 2x} = 1\] then 9/2x = 1 9 = 2x so x = 9/2?

OpenStudy (phi):

yes x= 4.5 hours (that is the fast machine) the other machine takes twice as long: 2x= 9 hours

OpenStudy (anonymous):

oh thank you so much ;3 i get it now